OpenStudy (anonymous):

Find the vertex F(x)=-2x^2+2x+6 the x and y intercepts line of symmetry max/min value

OpenStudy (mertsj):

x = -b/2a = -2/-4=1/2 y = -2(1/2)^2 +2(1/2) + 6 = -2(1/4)+1 + 6= -1/2 + 7 = 6 1/2 so Vertex is (1/2,6 1/2) Line of symmetry is a vertical line through the vertex so it is x = 1/2 The vertex represents the maximum/minimum. This parabola has a maximum since it is concave downward. The the maximum is 6 1/2 Find the y intercept by letting x = 0 so y = 6 and (0,6) is the y intercept. Find the x intercepts by letting y = 0 so -2x^2 + 2x + 6 = 0 or -2(x^2 - x - 3) = 0. Solve that using the quadratic equation.

OpenStudy (mertsj):

x = [1 + - sqrt(1-4(1)(-3)]/2 = [1 +- sqrt13]/2